OpenStudy (anonymous):

x^3 + 64 = 0. find the imaginary number by factoring??

OpenStudy (mertsj):

-4 is not an imaginary number so that cannot be the answer.

OpenStudy (mertsj):

Factor x^3+64

OpenStudy (mertsj):

I'll get you started: \[x^3+64=(x+4)(x^2 \]

OpenStudy (anonymous):

\[a^3+b^3=\left( a+b \right)\left( a^2-ab+b^2 \right)\]

OpenStudy (mertsj):

Now you finish the factoring.

OpenStudy (anonymous):

im a bit confused where you got x^3 + 64?? @Mertsj

OpenStudy (anonymous):

oh wow i needed to look at the problem im an idiot aha

OpenStudy (mertsj):

That is the problem you posted.

OpenStudy (mertsj):

You must have studied factoring and specifically the factoring of the sum of two cubes. Use that knowledge to factor x^3+64

OpenStudy (anonymous):

i must have i just do not remember. math is just very difficult for me

OpenStudy (mertsj):

Ok. Watch this: